OpenStudy (anonymous):

Salina currently has an account balance of $1,047.69. Her initial deposit on the account was $630 and it earned 3.9% simple interest. How long has Salina held the account? 17 years 26 years 10 years 43 years

OpenStudy (wolf1728):

Interest = 1,047.69 minus 630.00 which equals 417.69 and principal = 630.00 Simple Interest = rate * principal * # of time periods Simple Interest = .039 * 630.00 * # of time periods # of time periods = 417.69 / [.039 * 630] # of time periods = 417.69 / 24.57 # of time periods = 17 years

OpenStudy (texaschic101):

oh...I see my mistake....I was making 1047.69 be I instead of subtracting 630 from it. Your answer makes sense to me now

OpenStudy (anonymous):

Thanks a lot

OpenStudy (wolf1728):

u r welcome